OpenStudy (anonymous):

find the function rule x: -2, -1, 0, 1, 2 y: 9, 4, -1, -6, -22

OpenStudy (anonymous):

the function rule for x is x+1; so (-2)+1= -1, (-1)+1=0 and so on the function rule for y is x-5; so (-9)-5=4, (-4)-5=-1, (-1)-5=-6 etc.

OpenStudy (anonymous):

do you understand this?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

@godgavemeyou

OpenStudy (anonymous):

idk ... srry